# Protein Structure Tokenization via Geometric Byte Pair Encoding

### v1

@@ Abstract
Protein structure is central to biological function, and enabling multimodal protein models requires joint reasoning over sequence, structure, and function. A key barrier is the lack of principled protein structure tokenizers (PSTs): existing approaches fix token size or rely on continuous vector codebooks, limiting interpretability, multi-scale control, and transfer across architectures. We introduce GeoBPE, a geometry-grounded PST that transforms continuous, noisy, multi-scale backbone conformations into discrete “sentences” of geometry while enforcing global constraints. Analogous to byte-pair encoding, GeoBPE generates a hierarchical vocabulary of geometric primitives by iteratively (i) clustering Geo-Pair occurrences with k-medoids to yield a resolution-controllable vocabulary; (ii) quantizing each Geo-Pair to its closest medoid prototype; and (iii) reducing drift through differentiable inverse kinematics that optimizes boundary glue angles under an $\mathrm{SE}(3)$ end-frame loss. GeoBPE offers compression (>>10× reduction in bits-per-residue at similar distortion rate), data efficiency (>>10× less training data), and generalization (maintains test/train distortion ratio of 1.01.11.0-1.1). It is architecture-agnostic: (a) its hierarchical vocabulary provides a strong inductive bias for coarsening residue-level embeddings from large PLMs into motif- and protein-level representations, consistently outperforming leading PSTs across $12$ tasks and $24$ test splits; (b) paired with a transformer, GeoBPE supports unconditional backbone generation via language modeling; and (c) tokens align with CATH functional families and support expert-interpretable case studies, offering functional meaning absent in prior PSTs. Code is available at https://github.com/shiningsunnyday/PT-BPE/.
@@ 1 Introduction
Protein language models (PLMs) trained on large sequence databases capture evolutionary constraints (Rives et al., 2021) and support de novo sequence design (Lin et al., 2023b), but they do not explicitly model fold geometry and may underperform on tasks where function depends on structural interactions (Abramson et al., 2024; Gelman et al., 2025). In natural language processing, byte-pair encoding (BPE) constructs a vocabulary by iteratively merging the most frequent symbol pairs, producing a hierarchical representation of text (Larsson & Moffat, 2002). Despite BPE’s success on sequential data, there is no geometric analog that can encode and decode protein backbone conformations. The central difficulty is discretizing continuous, noisy structural variability while preserving global consistency. Because protein folds are organized into modular substructures (Petsko & Ringe, 2004), a protein structure tokenizer should (a) build a hierarchical vocabulary of structural motifs and (b) segment folds into hierarchical decompositions, producing symbolic and interpretable representations of backbone geometry.
@@ 2 Related Work
Protein Structural Alphabets. Structural alphabets approximate protein folds as successions of geometric motifs (Branden & Tooze, 2012). de Brevern et al. (2000) introduced 16 five–residue protein blocks from Protein Data Bank (PDB) structures, assigning fragments by RMSD. Later work showed that over 90% of residues can be covered by such alphabets (de Brevern et al., 2002) and analyzed their quality and specificity (de Brevern, 2005). Alphabet strings provide 1D encodings of 3D geometry, enabling the use of sequence alignment for fold analysis and prediction (Mahajan et al., 2015; Vetrivel et al., 2017). Camproux et al. (1999) proposed 12 building blocks via Hidden Markov Models (HMMs) and extended it to capture whole-protein conformational variability (Camproux et al., 2004). HMMs use inter-alpha-carbon distances within four residues as observed variables. Broader tertiary descriptors, such as inter-residue distances or moment invariants (Durairaj et al., 2020), capture non-contiguous context; Mackenzie (2016) found $\sim$600 motifs describe more than half of structural space ($39\cdot 10^{6}$ conformations), indicating variability collapses into limited modes. Such descriptors extend to protein-level retrieval and classification (Durairaj et al., 2020; Van Kempen et al., 2024; Barrio-Hernandez et al., 2023). GeoBPE builds on these insights by treating structural motifs as extensible primitives and dynamically adjusting alphabet size and token resolution, unlike fixed structural alphabets.
@@ 3.1 Notation & Preliminaries
The full internal representation thus contains $3N^{(\tau)}\!-\!1$ bond lengths, $3N^{(\tau)}\!-\!2$ bond angles, and $3N^{(\tau)}\!-\!3$ dihedrals and is invariant to any $g=(R,t)\in\mathrm{SE}(3)$.
Local Formulation (Bond–Residue). For residue $i$ we define the bond–residue as the ordered triple $(N_{i}\!-\!\mathrm{CA}_{i}),\quad(\mathrm{CA}_{i}\!-\!C_{i}),\quad(C_{i}\!-\!N_{i+1})$ together with its internal angles. For $i<N^{(\tau)}$ this includes the lengths $\ell^{N\!-\!CA}_{i},\ \ell^{CA\!-\!C}_{i},\ \ell^{C\!-\!N}_{i}$, the bond angles $\theta^{N\!CA\!C}_{i},\ \theta^{CA\!C\!N}_{i}$, and the peptide dihedral $\psi_{i}$ about $CA_{i}\!-\!C_{i}$. For $i=N^{(\tau)}$, it includes only bond lengths $\ell^{N\!-\!CA}_{N^{(\tau)}}$, $\ell^{CA\!-\!C}_{N^{(\tau)}},$ and angle $\theta^{N\!CA\!C}_{N^{(\tau)}}$ (the (CN)(C\!-\!N) bond, $\theta^{CA\!C\!N}_{N^{(\tau)}}$, $\theta^{C\!NCA}_{N^{(\tau)}}$, and ${\{\psi,\omega,\phi\}}_{{N^{(\tau)}}}$ dihedrals are absent).
@@
Motif Formulation. A bond–residue motif $\mathcal{M}_{p:q}$ is a contiguous block of bond–residues $i=p,\dots,q$ $(1\leq p\leq q\leq N^{(\tau)})$. Its internal parameters comprise $\big\{\ell^{N\!-\!CA}_{i},\ \ell^{CA\!-\!C}_{i},\ \ell^{C\!-\!N}_{i},\ \theta^{N\!CA\!C}_{i},\ \theta^{CA\!C\!N}_{i},\ \psi_{i}\big\}_{i=p}^{q}$ (with the obvious omissions at $i=q=N^{(\tau)}$) together with the internal glue angles $\{\Gamma_{i}\}_{i=p}^{q-1}$ that connect consecutive bond–residues inside the motif. The joined motif $\mathcal{M}_{p:r}$ requires the union of the internal parameters of $\mathcal{M}_{p:q}$ and $\mathcal{M}_{q+1:r}$, plus the external glue angles $\Gamma_{q}$ connecting the last and first bond-residues of $\mathcal{M}_{p:q}$ and $\mathcal{M}_{q+1:r}$.
@@
Per-Link Transform. Define the transform between consecutive residue frames $G_{i}\coloneqq F_{i+1}\,F_{i}^{-1}\ \in\ \mathrm{SE}(3).$ By construction, $G_{i}$ is a deterministic function of the internal coordinates local to the link $i\!\to\!i{+}1$, namely $G_{i}=g\!\big(\ell^{CA\!-\!C}_{i},\ \ell^{C\!-\!N}_{i},\ \ell^{N\!-\!CA}_{i+1},\ \theta^{N\!CA\!C}_{i},\ \theta^{CA\!C\!N}_{i},\ \theta^{C\!N\!CA}_{i},\ \psi_{i},\ \omega_{i},\ \phi_{i+1}\big)$, and, in particular, depends on the glue set $\Gamma_{i}=\{\theta^{C\!N\!CA}_{i},\ \phi_{i+1},\ \omega_{i}\}$.
@@ 3.2 GeoBPE Algorithm
GeoBPE (Algo. 1) is organized around four components: (1) clustering motif occurrences into representative structural modes, (2) maintaining an ordered map to track frequent Geo-Pairs, (3) adaptively quantizing Geo-Pairs with controllable resolution, and (4) applying rigid-body refinement to enforce global geometric consistency.
(1) Extracting Dominant Modes from a Set of Motif Occurrences. The core subroutine invoked by GeoBPE is Algo. 6, which clusters a set of length $L$ raw backbone fragments into $K$ representative prototypes. This induces a hard quantization of the fragment space, since every possible occurrence is assigned to exactly one prototype. Because RMSD defines a metric over fragments, the clustering yields a Voronoi partition of this space. Importantly, the medoids are themselves observed fragments, so each quantized symbol retains a concrete structural interpretation: it represents the closest empirically observed conformation, providing a denoised approximation of local variability. Each time we quantize, we substitute every non‑medoid occurrence by its assigned medoid, replacing all internal angles with the medoid’s internal angles, making it an exact copy of that medoid (same length and per‑position angles).
(2) Constructing a Structural Motif Alphabet. GeoBPE begins by quantizing all bond-residues and glue angles (Algo. 16) and building an ordered map (for dual-use as a priority queue, keyed by (occurrence count, grounding key)) of discrete geo-pair grounding keys to occurrences (Algo. 19). In each iteration (Algo. 9), GeoBPE pops the most frequent motif-pair key, runs Algo. 6 on all mapped occurrences, quantizes the occurrences, runs rigid-body refinement, and updates the dictionary to account for the new quantized backbone states.
@@
0: Backbones $\{t^{(1)},\dots,t^{(T)}\}$ with lengths $N^{(\tau)}$; residue codebook sizes $(K_{3},K_{2})$; glue-IK weights $(w_{R},w_{t})$; maximum merge iterations $S_{\max}$.
@@
2:Per-residue initialization (Algo. 16). Cluster interior and terminal bond–residues via RMSD_Partition to obtain codebooks $\mathcal{A}_{3},\mathcal{A}_{2}$; overwrite each residue’s internals by its assigned prototype. Set the initial segmentation for each backbone:
@@
3:Global glue refinement (Algo. 10). Optimize all boundary glues $\Gamma_{i}=\{\theta^{C\!N\!CA}_{i},\omega_{i},\phi_{i+1}\}$ via differentiable FK with $(w_{R},w_{t})$; snap each to the nearest bin center using $Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}$.
4:Build the priority-ordered geo-pair map (Algo. 19). Using the frontier leaves of each $\mathcal{F}^{(\tau)}$ (equivalently, $\mathcal{P}^{(\tau)}$), construct the occurrence sets $\mathcal{O}(\kappa)$ and insert:
@@
9:return $\mathcal{V}$, $\{\mathcal{P}^{(\tau)}\}$, $\{\mathcal{F}^{(\tau)}\}$, and $\mathcal{D}$.
@@
with forward kinematics $\widehat{F}_{k_{u}}\;=\;F^{\star}_{i_{u}-1}\,G^{\mathrm{new}}_{i_{u}-1}\,T^{\mathrm{med}}_{u},\qquad F^{\star}_{k_{u}}\;=\;F^{\star}_{i_{u}-1}\,G^{\mathrm{orig}}_{i_{u}-1}\,T^{\mathrm{occ}}_{u}.$ When quantizing many motifs on the same backbone, performing this optimization each time can become computationally prohibitive. Instead, we adopt a global (batch) alternative which treats all gluing DOFs as parameters, with a global end-frame fitting loss. This provides maximum flexibility in drift compensation. The algorithmic details are in Algo. 17 and 10.
Transferring Hierarchical Inductive Biases. GeoBPE can adapt the receptive field of a base feature extractor $\Theta$ to that of the whole structure, supporting fine-grained embeddings for localized predictions (e.g., active site prediction), coarse-grained embeddings for global predictions (e.g., fold classification), connecting both scales through recursive aggregation. Algo. 1 emits $\{\mathcal{F}^{(\tau)}\}$ as a forest, where leaf nodes represent residues and parent nodes represent motifs. Leaf nodes are initialized with pretrained features, then embeddings propagate up along the parent-child relations of $F^{(\tau)}$ until the forest roots; a final step aggregates the forest roots into a protein-level contextual embedding; then they are propagated down until the leaf nodes; the final embeddings are output as vocabulary-induced features. See Algo. 13 for details.

### v2

# arXiv:2511.11758v2[q-bio.QM] 01 Mar 2026

@@ Abstract
Protein structure is central to biological function, and enabling multimodal protein models requires joint reasoning over sequence, structure, and function. A key barrier is the lack of principled protein structure tokenizers (PSTs): existing approaches fix token size or rely on continuous vector codebooks, limiting interpretability, multi-scale control, and transfer across architectures. We introduce GeoBPE, a geometry-grounded PST that transforms continuous, noisy, multi-scale backbone conformations into discrete “sentences” of geometry while enforcing global constraints. Analogous to byte-pair encoding, GeoBPE generates a hierarchical vocabulary of geometric primitives by iteratively (i) clustering Geo-Pair occurrences with k-medoids to yield a resolution-controllable vocabulary; (ii) quantizing each Geo-Pair to its closest medoid prototype; and (iii) reducing drift through differentiable inverse kinematics that optimizes boundary glue angles under an $\mathrm{SE}(3)$ end-frame loss. GeoBPE offers compression (>>10× reduction in bits-per-residue at similar distortion rate), data efficiency (>>10× less training data), and generalization (maintains test/train distortion ratio of 1.01.11.0-1.1). It is architecture-agnostic: (a) its hierarchical vocabulary provides a strong inductive bias for coarsening residue-level embeddings from large PLMs into motif- and protein-level representations, consistently outperforming leading PSTs across $12$ tasks and $24$ test splits; (b) paired with a transformer, GeoBPE supports unconditional backbone generation via language modeling; and (c) tokens align with CATH functional families and support expert-interpretable case studies, offering functional meaning absent in prior PSTs. Code is available at https://github.com/shiningsunnyday/PT-BPE.
@@ 1 Introduction
Protein language models (PLMs) trained on large sequence databases capture evolutionary constraints (Rives et al., 2021) and support de novo sequence design (Lin et al., 2023b), but they do not explicitly model fold geometry and may underperform on tasks where function depends on structural interactions (Abramson et al., 2024; Gelman et al., 2025). In natural language processing, byte-pair encoding (BPE) constructs a vocabulary by iteratively merging the most frequent symbol pairs, producing a hierarchical representation of text (Larsson and Moffat, 2002). Despite BPE’s success on sequential data, there is no geometric analog that can encode and decode protein backbone conformations. The central difficulty is discretizing continuous, noisy structural variability while preserving global consistency. Because protein folds are organized into modular substructures (Petsko and Ringe, 2004), a protein structure tokenizer should (a) build a hierarchical vocabulary of structural motifs and (b) segment folds into hierarchical decompositions, producing symbolic and interpretable representations of backbone geometry.
@@ 2 Related Work
Protein Structural Alphabets. Structural alphabets approximate protein folds as successions of geometric motifs (Branden and Tooze, 2012). de Brevern et al. (2000) introduced 16 five–residue protein blocks from Protein Data Bank (PDB) structures, assigning fragments by RMSD. Later work showed that over 90% of residues can be covered by such alphabets (de Brevern et al., 2002) and analyzed their quality and specificity (de Brevern, 2005). Alphabet strings provide 1D encodings of 3D geometry, enabling the use of sequence alignment for fold analysis and prediction (Mahajan et al., 2015; Vetrivel et al., 2017). Camproux et al. (1999) proposed 12 building blocks via Hidden Markov Models (HMMs) and extended it to capture whole-protein conformational variability (Camproux et al., 2004). HMMs use inter-alpha-carbon distances within four residues as observed variables. Broader tertiary descriptors, such as inter-residue distances or moment invariants (Durairaj et al., 2020), capture non-contiguous context; Mackenzie (2016) found $\sim$600 motifs describe more than half of structural space ($39\cdot 10^{6}$ conformations), indicating variability collapses into limited modes. Such descriptors extend to protein-level retrieval and classification (Durairaj et al., 2020; Van Kempen et al., 2024; Barrio-Hernandez et al., 2023). GeoBPE builds on these insights by treating structural motifs as extensible primitives and dynamically adjusting alphabet size and token resolution, unlike fixed structural alphabets.
@@ 3.1 Notation & Preliminaries
Figure 2: Toy backbone, with internal angles, glues $T_{1}$, motif $\mathcal{M}_{1:2}$ and per-link transform $G_{1}$.
We annotate these definitions in a toy $(N^{(\tau)}=2)$} example in Fig. 2 (top). The full internal representation thus contains $3N^{(\tau)}\!-\!1$ bond lengths, $3N^{(\tau)}\!-\!2$ bond angles, and $3N^{(\tau)}\!-\!3$ dihedrals and is invariant to any $(R,t)\in\mathrm{SE}(3)$.
Local Formulation (Bond–Residue). For residue $i$ we define the bond–residue as the ordered triple $(N_{i}-\mathrm{CA}_{i}),(\mathrm{CA}_{i}\!-\!C_{i}),(C_{i}\!-\!N_{i+1})$ together with its internal angles. For $i<N^{(\tau)}$ this includes the lengths $\ell^{N\!-\!CA}_{i},\ \ell^{CA\!-\!C}_{i},\ \ell^{C\!-\!N}_{i}$, the bond angles $\theta^{N\!CA\!C}_{i},\ \theta^{CA\!C\!N}_{i}$, and the peptide dihedral $\psi_{i}$ about $CA_{i}\!-\!C_{i}$. For $i=N^{(\tau)}$, it includes only bond lengths $\ell^{N\!-\!CA}_{N^{(\tau)}}$, $\ell^{CA\!-\!C}_{N^{(\tau)}},$ and angle $\theta^{N\!CA\!C}_{N^{(\tau)}}$ (the (CN)(C\!-\!N) bond, $\theta^{CA\!C\!N}_{N^{(\tau)}}$, $\theta^{C\!NCA}_{N^{(\tau)}}$, and ${\{\psi,\omega,\phi\}}_{{N^{(\tau)}}}$ dihedrals are absent).
@@
Motif Formulation. A bond–residue motif $\mathcal{M}_{p:q}$ is a contiguous block of bond–residues $i=p,\dots,q$ $(1\leq p\leq q\leq N^{(\tau)})$. Its internal parameter set is the union over the internal bond lengths and angles of bond-residues $p,\ldots,q$ together with the internal glue angles $\{\Gamma_{i}\}_{i=p}^{q-1}$ that connect consecutive bond–residues inside the motif. Given $q\leq r\leq N^{(\tau)}$, we obtain a Geo-Pair occurrence ($\mathcal{M}_{p:q},\Gamma_{q},\mathcal{M}_{q:r}$) from the internal parameters of $\mathcal{M}_{p:q}$ and $\mathcal{M}_{q+1:r}$, plus the external glue angles $\Gamma_{q}$ connecting the last and first bond-residues of $\mathcal{M}_{p:q}$ and $\mathcal{M}_{q+1:r}$.
@@
Per-Link Transform. Define the transform between consecutive residue frames $G_{i}\ \coloneqq\ F_{i+1}\,F_{i}^{-1}\ \in\ \mathrm{SE}(3).$ By construction, $G_{i}$ is a deterministic function of the internal coordinates local to the link $i\!\to\!i{+}1$, namely $G_{i}\;=\;g\!\big(\ell^{CA\!-\!C}_{i},\ \ell^{C\!-\!N}_{i},\ \ell^{N\!-\!CA}_{i+1},\ \theta^{N\!CA\!C}_{i},\ \theta^{CA\!C\!N}_{i},\ \theta^{C\!N\!CA}_{i},\ \psi_{i},\ \omega_{i},\ {\phi_{i}}\big)$, and, in particular, depends on the glue set $\Gamma_{i}=\{\theta^{C\!N\!CA}_{i},\ {\phi_{i}},\ \omega_{i}\}$, illustrated in Fig. 2 (bottom).
@@
Figure 3: (Top) GeoBPE tracks a Geo-Pair Encoding, a dictionary mapping Geo-Pair keys to occurrences at all times. Each step pops the most frequent Geo-Pair key, gathers the occurrences ( , , , , ,…) and fixes $K$ prototypes ( , , ) to add to $\mathcal{V}$. All occurrences are quantized to the closest prototype (e.g. $\rightarrow$ ). Glue angles ( , ) are optimized to correct for the drift introduced. (Bottom) Toy example with two backbones; we initialize residue-orientation modes using two prototypes ( , ), pop the frequent Geo-Pair ( ), quantize occurrences ( $\rightarrow$ ), and optimize glue angles.
Core GeoBPE notation. We define the main objects used throughout the algorithmic description. We index training backbones by $\tau=1,\dots,T$, writing the $\tau$-th backbone as $t^{(\tau)}$ with $N^{(\tau)}$ residues. A segmentation of $t^{(\tau)}$ into bond–residue motifs is the ordered tuple $\mathcal{P}^{(\tau)}=\big(\mathcal{M}^{(t_{\tau})}_{p_{1}{:}q_{1}},\ldots,\mathcal{M}^{(t_{\tau})}_{p_{M_{\tau}}{:}q_{M_{\tau}}}\big)$, with $1=p_{1}\leq q_{1}<p_{2}\leq\cdots\leq q_{M_{\tau}}=N^{(\tau)}$. The corresponding merge hierarchy $\mathcal{F}^{(\tau)}$ is a binary forest whose frontier leaves, in order, equal $\mathcal{P}^{(\tau)}$; each internal node represents a merged motif and stores its span [p:q][p{:}q].
Geo-pair keys and occurrences.
Given two adjacent motifs $(\mathcal{M}_{p{:}q},\mathcal{M}_{q+1{:}r})$ and their boundary glue $\Gamma_{q}$, we define a canonical, hashable geo-pair key $\kappa=\textsc{ComputeGeoKey}(\mathcal{M}_{p{:}q},\mathcal{M}_{q+1{:}r})$ (Alg. 22). For each key $\kappa$ we collect its occurrence set $\mathcal{O}(\kappa)$ consisting of all such adjacent motif pairs across the dataset.
Prototypes and vocabulary.
For a geo-pair key $\kappa$, GeoBPE clusters its occurrences and stores a small set of representative prototypes $\mathcal{A}_{\kappa}=\{\Pi^{(\kappa)}_{j}\}_{j=1}^{K_{|\kappa|}}$, where each $\Pi^{(\kappa)}_{j}$ is the internal-parameter tuple of a medoid occurrence and $K:\mathbb{Z}^{+}\setminus\{1\}\mapsto\mathbb{Z}^{+}$ controls how many prototypes by motif (bond) length. The vocabulary is the map $\mathcal{V}:\ \kappa\mapsto\mathcal{A}_{\kappa}$, initially containing residue-level codebooks $\mathcal{A}_{3},\mathcal{A}_{2}$ and growing as new geo-pair keys are introduced.
Geo-pair dictionary and priorities.
At any time GeoBPE maintains a priority-ordered dictionary $\mathcal{D}$ that maps each key $\kappa$ to its occurrence set $\mathcal{O}(\kappa)$. Keys are ordered by tuples
$\pi(\kappa)=\big(\rho(\kappa),-|\mathcal{O}(\kappa)|,\kappa\big),\quad\rho(\kappa)=\mathbf{1}[\kappa\notin\mathrm{dom}(\mathcal{V})],$
so that compressible keys with existing prototypes ($\rho=0$) are popped first, followed by a new key with the largest count $|\mathcal{O}(\kappa)|$ per iteration.
@@ 3.2 GeoBPE Algorithm
GeoBPE (Algo. 1, Fig. 3) is organized around four components: (1) clustering motif (individual bond-residues once at the start, Geo-Pairs every step thereafter) occurrences into representative structural prototypes, (2) maintaining an ordered map to track frequent Geo-Pairs, (3) adaptively hard-quantizing noisy Geo-Pairs to their assigned prototypes, and (4) applying rigid-body refinement to enforce global geometric consistency.
Components (1)-(4) are designed to answer three new key questions when re-interpreting BPE to work with continuous backbone geometry rather than discrete bytes: (a) how do we ground continuous backbone states to discrete keys for Geo-Pair counting, (b) how do we update the backbone states once we have popped the most frequent Geo-Pair, (c) how do we synchronize the Geo-Pair dictionary with how we updated the backbone states. Component (1) answers to (a), (3) & (4) to (b), and (2) to (c). The guiding question encompassing (a)-(c) is: what is the exact relationship between the Geo-Pair dictionary (needed for discrete BPE operations) and the continuous backbone states (which should both reflect new keys and preserve original fidelity)? GeoBPE implements a two-way connection through four stages: grounding, in which continuous motif states define discrete prototype keys; quantization, in which internal parameters are overwritten with those of the assigned prototypes; rigid-body refinement, in which backbone internal states self-correct to minimize global distortion; and synchronization, in which the Geo-Pair dictionary is re-synchronized to reflect the high-fidelity backbone states after quantization and refinement.
(1) Extracting Dominant Modes from a Set of Motif Occurrences. The core subroutine invoked by GeoBPE is Algo. 6, which clusters a set of length $L$ raw backbone fragments into $K$ representative prototypes. This induces a hard quantization of the fragment space, since every possible occurrence is assigned to exactly one prototype. Because RMSD defines a metric over fragments, the clustering yields a Voronoi partition of this space. Importantly, the medoids are themselves observed fragments, so each quantized symbol retains a concrete structural interpretation: it represents the closest empirically observed conformation, providing a denoised approximation of local variability. Each time we quantize, we substitute every non‑medoid occurrence by its assigned medoid, replacing all internal parameters with the medoid’s internal parameters, making it an exact copy of that medoid (same length and per‑position angles).
(2) Constructing a Structural Motif Alphabet. GeoBPE begins by quantizing all bond-residues and glue angles (Algo. 18) and building an ordered map $\mathcal{D}$ of discrete geo-pair grounding keys $\kappa$} to occurrences $O(\kappa)$} (Algo. 21, see Core GeoBPE notation). In each iteration (Algo. 9), GeoBPE pops the most frequent Geo-Pair key, runs Algo. 6 on mapped occurrences, quantizes the occurrences, runs rigid-body refinement, and updates $\mathcal{D}$ to account for the new quantized backbone states.
@@
0: Backbones $\{t^{(1)},\dots,t^{(T)}\}$ with lengths $N^{(\tau)}$; optional backbones $\{t^{(\xi)}\}$ to tokenize; residue codebook sizes $(K_{3},K_{2})$; glue-IK weights $(w_{R},w_{t})$; maximum merge iterations $S_{\max}$.
@@
2:Per-residue initialization (Algo. 18). Cluster interior and terminal bond–residues via RMSD_Partition to obtain codebooks $\mathcal{A}_{3},\mathcal{A}_{2}$; overwrite each residue’s internals by its assigned prototype. Set the initial segmentation for each backbone:
@@
3:Global glue refinement (Algo. 12). Optimize all boundary glues $\Gamma_{i}=\{\theta^{C\!N\!CA}_{i},\omega_{i},\phi_{i+1}\}$ via differentiable FK with $(w_{R},w_{t})$; snap each to the nearest bin center using $Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}$.
4:Build the priority-ordered geo-pair map (Algo. 21). Using the frontier leaves of each $\mathcal{F}^{(\tau)}$ (equivalently, $\mathcal{P}^{(\tau)}$), construct the occurrence sets $\mathcal{O}(\kappa)$ and insert:
@@
9:Tokenize new/unseen backbones (Algo. 10) for each $\xi$,
$(\mathcal{P}^{(\xi)},\mathcal{F}^{(\xi)})\leftarrow\textsc{Tokenize}\big(t^{(\xi)},\mathcal{A}_{3},\mathcal{A}_{2},\mathcal{V},\{Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}\}\big)$}
10:return $\mathcal{V}$, $\{\mathcal{P}^{(\tau)}\}$, $\{\mathcal{F}^{(\tau)}\}$, $\mathcal{D}$ and (if given) $\{\mathcal{P}^{(\xi)}\}$, $\{\mathcal{F}^{(\xi)}\}$
@@
with forward kinematics $\widehat{F}_{k_{u}}\;=\;F^{\star}_{i_{u}-1}\,G^{\mathrm{new}}_{i_{u}-1}\,T^{\mathrm{med}}_{u},\qquad F^{\star}_{k_{u}}\;=\;F^{\star}_{i_{u}-1}\,G^{\mathrm{orig}}_{i_{u}-1}\,T^{\mathrm{occ}}_{u}.$ When quantizing many motifs on the same backbone, performing this optimization each time can become computationally prohibitive. Instead, we adopt a global (batch) alternative which treats all gluing DOFs as parameters, with a global end-frame fitting loss. This provides maximum flexibility in drift compensation. The algorithmic details are in Algo. 19 and 12.
Transferring Hierarchical Inductive Biases. GeoBPE adapts the receptive field of a base residue-level feature extractor $\Theta$ to that of the whole structure, connecting multiple scales through recursive aggregation. Algo. 1 emits merge hierarchies $\mathcal{F}$ as a forest: leaf nodes represent residues and parent nodes represent motifs. The key insight is to use F{F} as a recursive computation tree. Leaf nodes are initialized with pretrained features, then embeddings propagate up along the parent-child relations of $\mathcal{F}$ until the forest roots (aligned with $\mathcal{P}$); a final step aggregates the forest roots into a protein-level contextual embedding; then they are propagated down until the leaf nodes. The final leaf nodes output residue-level embeddings induced by the hierarchical $\mathcal{V}$ and informed by multi-scale GeoBPE tokenization. These features support supervised learning on fine-grained residue-level tasks (e.g., active site prediction) and coarse-grained global predictions (e.g., fold classification). See Algo. 15 for details.